OpenStudy (anonymous):

Factor completely: 4x^2 + 25x + 6

OpenStudy (nathan917):

4X^2 +25x+6 = (x+6)*(4x+1)

OpenStudy (anonymous):

none of those are an option

OpenStudy (anonymous):

(4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1)

OpenStudy (nathan917):

Id go With A.

OpenStudy (nathan917):

4x^2 + 25x + 6 = 4x^2 + 24x + x + 6 = 4x(x + 6) + 1(x + 6) = (x + 6)(4x + 1) Is how i got it.

OpenStudy (anonymous):

thnx

OpenStudy (nathan917):

(x + 6)(4x + 1) x times 4x = 4x^2 x times 1 = x 6 times 4x = 24x 6 times 1 = 6 Now add like terms and you get 4x^2 + 25x + 6 And since that matches you original problem, the answer of (x + 6)(4x + 1) is correct. Overall Answer is (x + 6)(4x + 1)